# QuadSync: Quadrifocal Tensor Synchronization via Tucker Decomposition

### v1

@@ 2 Background
    $rank(\mathcal{F}^{n})=6$ and $\mathcal{F}^{n}$ has exactly $3$ positive and $3$ negative eigenvalues.
@@ 2.1 Quadrifocal Tensors
There are certain interesting characteristics of quadrifocal tensors that motivate our choice to use them for synchronization. Firstly, quadrifocal tensors capture the interactions of four views, and thus encode complex geometric information. The quadrifocal tensor includes triple and pairwise information too. [35] showed the effectiveness of using redundant information in trifocal tensors for better averaging of translations, and quadrifocal tensors stand to further improve results. Secondly, there may be more flexibility for estimating quadrifocal tensors, especially in homogeneous scenes. In line correspondence relationships of quadrifocal tensors, the lines in the image planes do not necessarily have to correspond to the same world line. The quadrifocal tensor can also be estimated from $6$ point correspondences. We refer to the following for detecting and matching lines across views [33, 40]. Thirdly, in the case of quadrifocal tensors all views are treated equally, unlike in the trifocal tensor case where there is a view treated specially. This might lead to a more stable algorithm in synchronization. Lastly, but not least, estimating a quadrifocal tensor requires sufficiently many inlier features across four views. This implicitly enforces consistency of the viewing graph through matched points and/or lines. Intentionally using quadrifocal tensors for synchronization may therefore help us to start synchronization from a cleaner graph, especially in the dense viewing graph case. We refer to [20, 18, 37, 46] for more on the properties and estimation of quadrifocal tensors. We also refer to [14, 29, 30, 36, 26] for more on solving minimal problems related to computer vision. We refer to [27, 26, 37, 1, 25] for results on minimal problems and the characterization of higher order tensors from an algebraic geometry perspective.
@@ 7.1 Line Projection Matrices
where $A^{i}$ indicates the $i$th row of the camera matrix $A$. Here, $\wedge$ is the wedge product, or the exterior product, between two vectors. The order of the basis is given by the order in the columnindicescolumnindices in the following explanation. This can be calculated specifically in the following way. First, let $rowindices=[2,3;~1,3;~1,2],\hskip 2.5ptcolumnindices=[1,2;~1,3;~1,4;~2,3;~2,4;~3,4]$. Then let $\mathcal{P}$ be the $3\times 6$ matrix, where $\mathcal{P}_{i,j}$ is calculated by the minors
@@
if $i=2$. When $A$ is a $3n\times 4$ matrix, then let $\mathcal{P}$ denote the $3n\times 6$ matrix of the stacked exterior squares. We refer to [12] for a more detailed description of the exterior algebra of a vector space. Note that the matrix $\mathcal{P}$ is also the line projection matrices, which can project 3D world lines in plucker coordinates (which are in $\mathbb{R}^{6}$) to lines in the image plane. We refer to [20] for more details on the role of line projection matrices in computer vision.
@@ 8.1 Block Trifocal Tensor in Collinear Cases
Since it has been shown in the proof for Theorem 3.1 that the stacked camera matrices will only exhibit a low rank if all cameras lie at the same point, it is the same case for the low ranks in the flattenings of the second and third mode. We only need to show the rank drop in the first mode. Note that $\mathcal{T}^{n}_{(1)}=\mathcal{P}(\mathcal{G}_{T})_{(1)}(C\otimes C)^{T}$, we just need to show that $\mathcal{P}\in\mathbb{R}^{3n\times 6}$ has rank $5$. Since for individual line projection matrices $\mathcal{P}_{i}$, we have $\mathcal{P}_{i}\mathcal{L}=0$ only when $\mathcal{L}$ is a line that passes through the camera center. Since we have a collinear set of cameras, there is only one line that passes through all of the camera centers, hence the kernel of $\mathcal{P}$ has dimension 1. ∎
@@
    when $x_{3}y_{4}-x_{4}y_{3}\not=0$. Thus, for generic $X$, $\text{rank}(\mathcal{G}_{S})=2$, so that $\text{rank}(C\mathcal{G}_{S}C^{T})=2$, proving the projection rank for $M_{1}$ to be $2$. This has also been checked numerically and symbolically.
@@
    For $\mathcal{Q}^{n}_{iiii}$ diagonal blocks, all four cameras will be the same in the definition for quadrifocal tensors from camera matrices. Since $P_{i}$ only has $3$ rows, there must be a repeating row in the determinant, meaning that $\mathcal{Q}^{n}_{iiii}=0$ in $\mathbb{R}^{3\times 3\times 3\times 3}$.
@@
    Suppose that we now have $3$ overlapping indices, such that we are considering the blocks $\mathcal{Q}^{n}_{iiij},\mathcal{Q}^{n}_{iiji},\mathcal{Q}^{n}_{ijii},\mathcal{Q}^{n}_{jiii}$ for $i\not=j$. Then, among the four rows in the determinant, three must correspond to the same camera $i$. By [20], these entries will be the epipoles up to signs, where epipoles are the images of the camera center of view $i$ in view $j$.
@@
    Suppose that we now have $2$ overlapping indices, such that we are looking at the blocks $\mathcal{Q}^{n}_{iijk}$, $\mathcal{Q}^{n}_{ijik}$, $\mathcal{Q}^{n}_{ijki}$, $\mathcal{Q}^{n}_{jiik}$, $\mathcal{Q}^{n}_{jiki}$, $\mathcal{Q}^{n}_{jkii}$, $\mathcal{Q}^{n}_{iikj}$, $\mathcal{Q}^{n}_{ikij}$, $\mathcal{Q}^{n}_{ikji}$, $\mathcal{Q}^{n}_{kiij}$, $\mathcal{Q}^{n}_{kiji}$, $\mathcal{Q}^{n}_{kjii}$ for $i\not=j,i\not=k,j\not=k$. Then, among the four rows in the determinant, two correspond to the same camera $i$. Again by [20], these entries will be the elements of the trifocal tensor up to signs.
@@
    Suppose that we now have $2$ overlapping indices, and the other two are also the same, such that we are looking at the blocks, $\mathcal{Q}^{n}_{iijj},\mathcal{Q}^{n}_{ijij},\mathcal{Q}^{n}_{ijji},\mathcal{Q}^{n}_{jiij},\mathcal{Q}^{n}_{jiji},\mathcal{Q}^{n}_{jjii}$ for $i\not=j$. Then, among the four rows in the determinant, two correspond to the same camera $i$, and the other two correspond to the same camera $j$. Again by [20], these entries will be the elements of the fundamental matrices up to signs.
@@
  • $\lambda_{ijk\ell}=c$ if exactly two of $i,j,k,\ell$ equal $1$;
  • $\lambda_{ijk\ell}=c^{2}$ if exactly one of $i,j,k$ equals $1$;
  • $\lambda_{ijk\ell}=c^{3}$ if none of $i,j,k,\ell$ equal $1$ and the indices are not all the same.
@@
We let $\mathcal{Q}^{n}_{(1)}$ and $(\lambda\odot_{b}\mathcal{Q}^{n})_{(1)}$ denote the mode-$1$ matrix flattenings in $\mathbb{R}^{3n\times 27n^{3}}$ of the block quadrifocal tensor and its scaled variant, where rows correspond to the first mode of the tensors. Invoking Theorem 3.1 alongside our standing assumptions shows that both matrices have rank $4$. Consequently, every $5\times 5$ minor of these matrices must be zero.
@@
We now see that $\lambda$-entries with two $1$-indices agree. Indeed, taking $i=j$ above gives $\lambda_{\pi_{1}(i1i1)}=\lambda_{\pi_{2}(ii11)}$ for all $\pi_{1}$ and $\pi_{2}$ that fix $(ii11)$ and $(i1i1)$ respectively. So $\lambda_{ii11}=\lambda_{\pi(ii11)}$ for all $\pi$. Taking $i\neq j$ gives $\lambda_{ii11}=\lambda_{\pi(i1j1)}=\lambda_{jj11}$ for all $\pi$. Together, there exists $c\in\mathbb{R}^{*}$ such that $c=\lambda_{\pi(ij11)}$ for all $i,j\in\{2,\dots,n\}$ and permutations $\pi$.
@@
where asterisks denote corresponding entries in $\mathcal{Q}^{n}_{(1)}$. As a polynomial in $c$ and $\lambda_{ijk1}$, the determinant of (14) is a scalar multiple of $c(c^{2}-\lambda_{ijk1})$. This is because the polynomial has degree $\leq 3$, if $c=0$ then the bottom two rows of (14) are linearly dependent, and if $c^{2}=\lambda_{ijk1}$ then (14) is a $5\times 5$ submatrix of $\mathcal{Q}^{n}_{(1)}$ with row and column operations performed. The scale is a polynomial in $P_{1},P_{i},P_{j},P_{k}$. It is Zariski-generically nonzero if we exhibit one instance of camera matrices such that the determinant of (13) does not vanish for all $c,\lambda_{ijk1}$. Further, it suffices to find an instance where $i=j=k$, as all other cases specialize to this. Computational verification with a random numerical instance of $P_{1},P_{i}$ proves the non-vanishing. It follows that $c^{2}=\lambda_{ijk1}$. Appealing to symmetry like before, $c^{2}=\lambda_{\pi(ijk1)}$ for all $i,j,k\in\{2,\dots,n\}$ and permutations $\pi$. Summarizing, all $\lambda$-entries with a single $1$-index equal $c^{2}$.
@@
The determinant of (15) is $c(c^{3}-\lambda_{ijk\ell})$ multiplied by a polynomial in $P_{1},P_{i},P_{j},P_{k},P_{\ell}$. The most specialized case is $i=j=k$. Computer verification with a random numerical instance proves the polynomial is not identically zero. We deduce that $c^{3}=\lambda_{ijk\ell}$. By symmetry, $c^{3}=\lambda_{\pi(ijk\ell)}$ for all $i,j,k,\ell\in\{2,\dots,n\}$ with $i,\ell$ distinct and all permutations $\pi$. In other words, $\lambda$-entries with no $1$-indices and non-identical indices equal $c^{3}$.
@@
    Generically, this matrix $\mathcal{T}^{n}_{1}$ will clearly have rank $4$.
@@
    This matrix generically will clearly have rank $3$. Denote $R_{i}$ as the $i$-th row of $\mathcal{T}^{n}_{2}$. We can observe that the fourth row is a linear combination of rows 1,2,3, where $y_{4}R_{4}=y_{1}R_{1}+y_{2}R_{2}+y_{3}R_{3}$. However, $R_{1},R_{2},R_{3}$ must be linearly independent in general.
@@ 9.1 QuadSync
For the flop complexity, we first account for the initialization. For the initialization, $C$ is calculated via HOSVD of a matrix of size $3n\times 27n^{3}$ where the target rank is $4$. This costs costs $O(4\times 81\times n^{4})=O(n^{4})$ flops. For the running iterations, to calculate $C$, one has to multiply a matrix of size $4\times 27n^{3}$ with a diagonal matrix of size $27n^{3}\times 27n^{3}$, then take the product and multiply it with another matrix of size $27n^{3}\times 4$. This multiplication of three matrix will have to be done for a total of $n$ times for each update of the stacked camera matrices $C$. Thus, the flop complexity for calculating the updates for $C$ is $O(n^{4})$. Calculating $\Lambda_{Q},~W$ also requires $O(|\Omega|)=O(n^{4})$ flops. Calculating $B,~\Gamma_{i}$ requires $O(n)$ flops. The overall flop complexity is $O(n^{4})$.

### v2

@@ Abstract
Complementary code can be found at https://github.com/dmiao153/QuadSync.
@@ 2 Background
    $rank(\mathcal{F}^{n})=6$ and $\mathcal{F}^{n}$ has exactly 33 positive and 33 negative eigenvalues.
@@ 2.1 Quadrifocal Tensors
There are certain interesting characteristics of quadrifocal tensors that motivate our choice to use them for synchronization. Firstly, quadrifocal tensors capture the interactions of four views, and thus encode complex geometric information. The quadrifocal tensor includes triple and pairwise information too. [35] showed the effectiveness of using redundant information in trifocal tensors for better averaging of translations, and quadrifocal tensors stand to further improve results. Secondly, there may be more flexibility for estimating quadrifocal tensors, especially in homogeneous scenes. In line correspondence relationships of quadrifocal tensors, the lines in the image planes do not necessarily have to correspond to the same world line. The quadrifocal tensor can also be estimated from 66 point correspondences. We refer to the following for detecting and matching lines across views [33, 40]. Thirdly, in the case of quadrifocal tensors all views are treated equally, unlike in the trifocal tensor case where there is a view treated specially. This might lead to a more stable algorithm in synchronization. Lastly, but not least, estimating a quadrifocal tensor requires sufficiently many inlier features across four views. This implicitly enforces consistency of the viewing graph through matched points and/or lines. Intentionally using quadrifocal tensors for synchronization may therefore help us to start synchronization from a cleaner graph, especially in the dense viewing graph case. We refer to [20, 18, 37, 46] for more on the properties and estimation of quadrifocal tensors. We also refer to [14, 29, 30, 36, 26] for more on solving minimal problems related to computer vision. We refer to [27, 26, 37, 1, 25] for results on minimal problems and the characterization of higher order tensors from an algebraic geometry perspective.
@@ 7.1 Line Projection Matrices
where $A^{i}$ indicates the $i$th row of the camera matrix AA. Here, $\wedge$ is the wedge product, or the exterior product, between two vectors. The order of the basis is given by the order in the columnindicescolumnindices in the following explanation. This can be calculated specifically in the following way. First, let $rowindices=[2,3;~1,3;~1,2],\hskip 2.5ptcolumnindices=[1,2;~1,3;~1,4;~2,3;~2,4;~3,4]$. Then let $\mathcal{P}$ be the $3\times 6$ matrix, where $\mathcal{P}_{i,j}$ is calculated by the minors
@@
if $i=2$. When AA is a $3n\times 4$ matrix, then let $\mathcal{P}$ denote the $3n\times 6$ matrix of the stacked exterior squares. We refer to [12] for a more detailed description of the exterior algebra of a vector space. Note that the matrix $\mathcal{P}$ is also the line projection matrices, which can project 3D world lines in plucker coordinates (which are in $\mathbb{R}^{6}$) to lines in the image plane. We refer to [20] for more details on the role of line projection matrices in computer vision.
@@ 8.1 Block Trifocal Tensor in Collinear Cases
Since it has been shown in the proof for Theorem 3.1 that the stacked camera matrices will only exhibit a low rank if all cameras lie at the same point, it is the same case for the low ranks in the flattenings of the second and third mode. We only need to show the rank drop in the first mode. Note that $\mathcal{T}^{n}_{(1)}=\mathcal{P}(\mathcal{G}_{T})_{(1)}(C\otimes C)^{T}$, we just need to show that $\mathcal{P}\in\mathbb{R}^{3n\times 6}$ has rank 55. Since for individual line projection matrices $\mathcal{P}_{i}$, we have $\mathcal{P}_{i}\mathcal{L}=0$ only when $\mathcal{L}$ is a line that passes through the camera center. Since we have a collinear set of cameras, there is only one line that passes through all of the camera centers, hence the kernel of $\mathcal{P}$ has dimension 1. ∎
@@
    when $x_{3}y_{4}-x_{4}y_{3}\not=0$. Thus, for generic $X$, $\text{rank}(\mathcal{G}_{S})=2$, so that $\text{rank}(C\mathcal{G}_{S}C^{T})=2$, proving the projection rank for $M_{1}$ to be 22. This has also been checked numerically and symbolically.
@@
    For $\mathcal{Q}^{n}_{iiii}$ diagonal blocks, all four cameras will be the same in the definition for quadrifocal tensors from camera matrices. Since $P_{i}$ only has 33 rows, there must be a repeating row in the determinant, meaning that $\mathcal{Q}^{n}_{iiii}=0$ in $\mathbb{R}^{3\times 3\times 3\times 3}$.
@@
    Suppose that we now have 33 overlapping indices, such that we are considering the blocks $\mathcal{Q}^{n}_{iiij},\mathcal{Q}^{n}_{iiji},\mathcal{Q}^{n}_{ijii},\mathcal{Q}^{n}_{jiii}$ for $i\not=j$. Then, among the four rows in the determinant, three must correspond to the same camera ii. By [20], these entries will be the epipoles up to signs, where epipoles are the images of the camera center of view $i$ in view $j$.
@@
    Suppose that we now have 22 overlapping indices, such that we are looking at the blocks $\mathcal{Q}^{n}_{iijk}$, $\mathcal{Q}^{n}_{ijik}$, $\mathcal{Q}^{n}_{ijki}$, $\mathcal{Q}^{n}_{jiik}$, $\mathcal{Q}^{n}_{jiki}$, $\mathcal{Q}^{n}_{jkii}$, $\mathcal{Q}^{n}_{iikj}$, $\mathcal{Q}^{n}_{ikij}$, $\mathcal{Q}^{n}_{ikji}$, $\mathcal{Q}^{n}_{kiij}$, $\mathcal{Q}^{n}_{kiji}$, $\mathcal{Q}^{n}_{kjii}$ for $i\not=j,i\not=k,j\not=k$. Then, among the four rows in the determinant, two correspond to the same camera ii. Again by [20], these entries will be the elements of the trifocal tensor up to signs.
@@
    Suppose that we now have 22 overlapping indices, and the other two are also the same, such that we are looking at the blocks, $\mathcal{Q}^{n}_{iijj},\mathcal{Q}^{n}_{ijij},\mathcal{Q}^{n}_{ijji},\mathcal{Q}^{n}_{jiij},\mathcal{Q}^{n}_{jiji},\mathcal{Q}^{n}_{jjii}$ for $i\not=j$. Then, among the four rows in the determinant, two correspond to the same camera $i$, and the other two correspond to the same camera $j$. Again by [20], these entries will be the elements of the fundamental matrices up to signs.
@@
  • $\lambda_{ijk\ell}=c$ if exactly two of $i,j,k,\ell$ equal 11;
  • $\lambda_{ijk\ell}=c^{2}$ if exactly one of $i,j,k$ equals 11;
  • $\lambda_{ijk\ell}=c^{3}$ if none of $i,j,k,\ell$ equal 11 and the indices are not all the same.
@@
We let $\mathcal{Q}^{n}_{(1)}$ and $(\lambda\odot_{b}\mathcal{Q}^{n})_{(1)}$ denote the mode-11 matrix flattenings in $\mathbb{R}^{3n\times 27n^{3}}$ of the block quadrifocal tensor and its scaled variant, where rows correspond to the first mode of the tensors. Invoking Theorem 3.1 alongside our standing assumptions shows that both matrices have rank 44. Consequently, every $5\times 5$ minor of these matrices must be zero.
@@
We now see that $\lambda$-entries with two 11-indices agree. Indeed, taking $i=j$ above gives $\lambda_{\pi_{1}(i1i1)}=\lambda_{\pi_{2}(ii11)}$ for all $\pi_{1}$ and $\pi_{2}$ that fix $(ii11)$ and $(i1i1)$ respectively. So $\lambda_{ii11}=\lambda_{\pi(ii11)}$ for all $\pi$. Taking $i\neq j$ gives $\lambda_{ii11}=\lambda_{\pi(i1j1)}=\lambda_{jj11}$ for all $\pi$. Together, there exists $c\in\mathbb{R}^{*}$ such that $c=\lambda_{\pi(ij11)}$ for all $i,j\in\{2,\dots,n\}$ and permutations $\pi$.
@@
where asterisks denote corresponding entries in $\mathcal{Q}^{n}_{(1)}$. As a polynomial in $c$ and $\lambda_{ijk1}$, the determinant of (14) is a scalar multiple of $c(c^{2}-\lambda_{ijk1})$. This is because the polynomial has degree $\leq 3$, if $c=0$ then the bottom two rows of (14) are linearly dependent, and if $c^{2}=\lambda_{ijk1}$ then (14) is a $5\times 5$ submatrix of $\mathcal{Q}^{n}_{(1)}$ with row and column operations performed. The scale is a polynomial in $P_{1},P_{i},P_{j},P_{k}$. It is Zariski-generically nonzero if we exhibit one instance of camera matrices such that the determinant of (13) does not vanish for all $c,\lambda_{ijk1}$. Further, it suffices to find an instance where $i=j=k$, as all other cases specialize to this. Computational verification with a random numerical instance of $P_{1},P_{i}$ proves the non-vanishing. It follows that $c^{2}=\lambda_{ijk1}$. Appealing to symmetry like before, $c^{2}=\lambda_{\pi(ijk1)}$ for all $i,j,k\in\{2,\dots,n\}$ and permutations $\pi$. Summarizing, all $\lambda$-entries with a single 11-index equal $c^{2}$.
@@
The determinant of (15) is $c(c^{3}-\lambda_{ijk\ell})$ multiplied by a polynomial in $P_{1},P_{i},P_{j},P_{k},P_{\ell}$. The most specialized case is $i=j=k$. Computer verification with a random numerical instance proves the polynomial is not identically zero. We deduce that $c^{3}=\lambda_{ijk\ell}$. By symmetry, $c^{3}=\lambda_{\pi(ijk\ell)}$ for all $i,j,k,\ell\in\{2,\dots,n\}$ with $i,\ell$ distinct and all permutations $\pi$. In other words, $\lambda$-entries with no 11-indices and non-identical indices equal $c^{3}$.
@@
    Generically, this matrix $\mathcal{T}^{n}_{1}$ will clearly have rank 44.
@@
    This matrix generically will clearly have rank 33. Denote $R_{i}$ as the $i$-th row of $\mathcal{T}^{n}_{2}$. We can observe that the fourth row is a linear combination of rows 1,2,3, where $y_{4}R_{4}=y_{1}R_{1}+y_{2}R_{2}+y_{3}R_{3}$. However, $R_{1},R_{2},R_{3}$ must be linearly independent in general.
@@ 9.1 QuadSync
For the flop complexity, we first account for the initialization. For the initialization, $C$ is calculated via HOSVD of a matrix of size $3n\times 27n^{3}$ where the target rank is 44. This costs costs $O(4\times 81\times n^{4})=O(n^{4})$ flops. For the running iterations, to calculate $C$, one has to multiply a matrix of size $4\times 27n^{3}$ with a diagonal matrix of size $27n^{3}\times 27n^{3}$, then take the product and multiply it with another matrix of size $27n^{3}\times 4$. This multiplication of three matrix will have to be done for a total of $n$ times for each update of the stacked camera matrices $C$. Thus, the flop complexity for calculating the updates for $C$ is $O(n^{4})$. Calculating $\Lambda_{Q},~W$ also requires $O(|\Omega|)=O(n^{4})$ flops. Calculating $B,~\Gamma_{i}$ requires $O(n)$ flops. The overall flop complexity is $O(n^{4})$.